# TPIA: Towards Target-specific Prompt Injection Attack against Code-oriented LLMs

### v5

@@ III-A Code Completion Task
Code completion is an essential task for Code LLMs. The input to the models may include code snippets, comments, pseudocode, and natural language requests. The output should consist of executable code blocks that meet the requirements specified by the input. In this paper, we consider a general task format. Specifically, the input code block is defined as $\bm{X}=\{x_{1},x_{2},x_{3},...,x_{m}\}$, where $x_{i}\in\bm{X}$ can be any token, and $\bm{X}$ as a whole adheres to the code format requirements. For instance, natural language requests can only be included in the code comments. The output is the completed code block $\bm{Y}=\{x_{1},x_{2},x_{3},...,x_{m},x_{m+1},x_{m+2},...,x_{n}\}$, which should be logically executable and functionally compliant with the requirements. We denote a Code LLM with weights $\bm{\theta}$ as $\Theta$, and the code completion task is simplified as $\bm{Y}=\Theta(\bm{X})$.
III-B Target-specific Attack
In the code domain, target-specific attack refers to the scenario where the attacker makes the Code LLM generate unsafe code for specific targets. Generally, the attacker’s targets can be described as a tuple of two continuous token sequences $(\bm{Y_{T}},\bm{T})$, where $\bm{Y_{T}}=\{x^{t}_{1},x^{t}_{2},...,x^{t}_{j}\}$ is a subsequence within $\bm{Y}$, and $\bm{T}=\{t_{1},t_{2},t_{3},...,t_{k}\}$ is what the attacker aims to insert after $\bm{Y_{T}}$ to create malicious functionality. To achieve the targets, the attacker needs to design a perturbation $\bm{P}$, represented as $\bm{P}=F(\bm{Y_{T}},\bm{T})$, and insert it into the input. The function $F$ can be implemented as manually creating the perturbation or by an optimization-based method as proposed in this paper. It is worth noting that, unlike backdoor attacks, the connection between $\bm{Y_{T}}$ and $\bm{T}$ is not built during the training process. To conclude, for a successful target-specific attack, the model output received by the victim $\bm{Y_{v}}$ will be:
$\begin{split}Y_{v}&=\Theta(\bm{X}\oplus\bm{P})\\ &=\{x_{1},x_{2},...,x^{t}_{1},x^{t}_{2},...,x^{t}_{j},t_{1},t_{2},...,t_{k},...,x_{n}\},\end{split}$ (1)
where $\oplus$ means contacting the two token sequences.
Figure 2: The overall pipeline of our TPIA attack. It is executed with two attack steps. (1) Objective Transfer. The attacker constructs the optimization task based on the malicious objective and generates a concise non-functional perturbation (instead of functional code used in supply chain attacks), containing instructions to achieve the constructed objective. (2) Perturbation Injection. The malicious code containing the malicious perturbation will be submitted to the open-source platforms (e.g., GitHub [3] and StackOverflow [5]) and introduced to the victim’s code context via various paths, such as direct copy and installing packages. When the victim requests the Code LLM to perform code completion based on his code context, the perturbation will induce the model to respond to the malicious instructions at specific locations.
III-C Threat Model
Attack Scenarios. We consider common scenarios where users introduce external code by copying or installing pack- ages [22, 41]. The attacker generates a non-functional perturbation to circumvent detection methods (e.g., static analysis [33] and sandbox [9]) and inserts it into victim’s code context via the external code. When the user requests the Code LLM to perform code completion, the perturbation will induce the model to generate an unsafe output at the specified location. Besides introducing code to the victim’s local project, attackers can also upload their codes containing perturbation to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded as a knowledge base to enhance the performance of Code LLMs [13] via retrieval-augmented generation [27]. Perturbation in the knowledge base can also appear in the context of the Code LLM’s prompts.
Attacker’s Capabilities. We assume that the attacker can obtain the parameters of an assistive Code LLM to generate malicious perturbations. This assumption is consistent with many adversarial methods [63, 50, 24]. Specifically, the attacker ideally obtains the open-source model corresponding to the target application, or uses the most effective assistive model in gray-box and black-box scenarios. On the other hand, perturbations can be detected and removed. Therefore, the attacker needs to employ non-functional perturbations rather than the functional code commonly used in supply chain attacks [41], since the latter can be easily detected by existing methods (e.g., static analysis [33] and sandboxing [9]).
Attacker’s Goals. The attacker has a specific malicious objective, which can be represented as a tuple (target position, target code). This tuple follows the definition of target-specific attacks in Section III-B. Specifically, the target position can be represented as $\bm{Y_{T}}$, and the target code can be represented as $\bm{T}$. The attacker attempts to insert a piece of designed malicious code $\bm{T}$ after a segment of the victim’s normal code $\bm{Y_{T}}$, and cause the victim to receive abnormal output code $\bm{Y_{v}}$. To achieve this goal, the attacker applies the adversarial methods to compute a perturbation $\bm{P}$ from the Code LLM’s parameters $\bm{\theta}$, which is most likely to induce the Code LLM to complete $\bm{T}$ after $\bm{Y_{T}}$. It can be formulated as below:
$\begin{array}[]{c}\bm{P}=\text{Adv}(\bm{\theta},\bm{Y_{T}},\bm{T}).\end{array}$ (2)
Note that Equation (2) can be regarded as the ‘adversarial attack version’ of $F(\bm{Y_{T}},\bm{T})$ defined in Section III-B and the generated perturbation $\bm{P}$ should satisfy Equation (1).
In particular, the attacker can achieve these goals by inserting code snippets into specified positions, without needing to oppose the Code LLMs’ safety alignment. For example, when setting the target position to DES.new( and the target code snippet to key,DES.MODE_ECB), the generated perturbation can induce Code LLMs to use the easily cracked ECB mode for encryption; When setting the target position to TRG() (a shorthand for Target which can be replaced with any function name), the generated perturbation can induce Code LLMs to insert designed code snippet into the TRG function or after using it.
IV Methodology
IV-A Overview
We hereby briefly introduce our target-specific prompt injection attack (TPIA). In general, TPIA consists of two main stages: task construction and malicious perturbation generation, as shown in Figure 2. Specifically, in the first stage, we construct the input and target output based on the attacker-specified objective (Section IV-B). In the second stage, we generate the malicious non-functional perturbation to achieve the constructed objective (Section IV-C). The perturbation generation process in the second stage involves two optimizations: (1) forward reasoning enhancement aligns the optimization goals of adversarial computation with Code LLMs’ sampling reasoning process (Section IV-D); (2) keyword-based perturbation design reduces the token usage of the perturbation to enhance the stealthiness (Section IV-E). We describe their technical details in the following parts.
IV-B Task Construction
In the first stage, TPIA constructs specific code completion tasks based on the malicious objectives. The constructed tasks simulate the victim’s local code environment and are used to optimize the malicious perturbation, which is closely related to the final performance of TPIA. Each task can be divided into two parts: the input containing the target position, and the expected output containing the target code. This task is to be constructed so that our attack satisfies the following three properties to ensure effectiveness.
@@ IV-B Task Construction
Property 1: Feasible Malicious Intent. The code environment configured by the user must meet the basic requirements for the malicious objective, including the necessary packages and methods declarations. Therefore, we introduce the conditional code at the beginning of the input code.
Property 2: Accurate Code LLMs’ Response. The response of the Code LLMs should be the target position. As stated in Section III, the target code should be inserted after the target position to achieve the malicious objective. Hence, we place the position code at the end of the input, maintaining its continuity with the target code.
Property 3: Robust Malicious Perturbation. The attacker cannot anticipate user’s contextual code in advance, which is the inevitable noise for TPIA. Therefore, we design the context code and add irrelevant noise to enhance the robustness of the malicious perturbation. The context code is placed between the position code and conditional code.
@@
To better illustrate our strategy, we provide a straightforward example in Figure 3, where the malicious objective is to make the victim use a weak encryption mode, i.e., making the Code LLM insert Target Code (‘‘DES.MODE_ECB’’) after “cipher = DES.new(key,”. To achieve this, TPIA constructs an output tuple as (“cipher = DES.new(key,” , “DES.MODE_ECB)”. The other three task components are automatically generated with the following stages. Firstly, we declare the Crypto series methods that enable the encryption code to run normally in the conditional code. Secondly, we add “cipher = DES.new(key,” to the last line of the position code to ensure the continuity with the target output “DES.MODE_ECB)”. In this example we also consider the necessary parameter “key” for the encryption method, as well as the disturbance parameter “iv”, which is required by encryption methods like “CBC” but not needed by “ECB”. Finally, TPIA constructs the code unrelated to the task and uses it as the noise-like context code. Particularly, TPIA constructs context codes of multiple aspects for the same task to collaboratively enhance the robustness.
Figure 3: The example of task construction that induces Code LLM to use encryption mode with low security. In the entire demo file, TPIA simulates the parts related to the malicious objective and constructs the tasks accordingly.
IV-C Malicious Perturbation Generation
In the second stage, TPIA optimizes the perturbation through an adversarial gradient computation method such that it can most likely induce the Code LLMs to generate the target malicious code. Utilizing the task code constructed in the first stage and the probabilities output by the target model, TPIA calculates the gradient of each token in the perturbation at the embedding layer. Through a greedy search approach, the optimal perturbation is derived. The optimization process is illustrated in Algorithm 1, with detailed descriptions below.
Algorithm 1 Perturbation Tokens Optimization
@@ IV-C Malicious Perturbation Generation
1:  Input: Task Code $\bm{T_{s}}$, target code $\bm{T}$, initial perturbation $\bm{P}$ with adversarial parts $x_{\text{adv},1:l}$, iterations $k$
2:  Vocab set $V$, $\mathbf{e}\in V$
@@
5:        $\mathcal{L}=\text{get\_loss}(\bm{T_{s}},\bm{T},\bm{P},\bm{\theta})$
6:        $\mathcal{G}=\nabla_{x_{\text{adv},i}}\mathcal{L}$
7:        $\mathbf{x_{\text{candidates}}}=\text{Top-}k_{e^{\prime}_{i}}\text{min}({e^{\prime}_{i}}^{T}\mathcal{G}),e^{\prime}_{i}\in V$
@@
9:           $\bm{P}^{\prime}:x_{\text{adv},i}\to x_{\text{candidates}}^{j}$
10:           $\mathcal{L}^{\prime}=\text{get\_loss}(\bm{T_{s}},\bm{T},\bm{P^{\prime}},\bm{\theta})$
@@
16:  until perturbation did not change in the last iteration
@@
Firstly, TPIA sets the adversarial token count $l$ and randomly initializes the perturbation. Considering that our enhancement method adds static tokens to the perturbation (Section IV-E), here we denote the adversarial part of the perturbation as $x_{\text{adv}_{1:l}}$. After determining the input and output of the code completion task, TPIA starts to optimize the $x_{\text{adv}_{1:l}}$ token by token in each iter (Line 4). We design a loss function based on the probability of the Code LLM for generating the target output (Line 5), which can be formulated as follows:
$\mathcal{L}=$ $\sum_{i=1}^{|\bm{T}|}\log(1-p(t_{i}|\bm{T_{s}}\oplus\bm{P},t_{1},t_{2},\ldots,t_{i-1})),$ (3)
where $\bm{T_{s}}$ denotes the task code. We only consider the prediction probability of each token in the target code when designing the loss function since the success of TPIA is solely related to whether the target code is correctly generated.
Besides, the optimization objective for the $i$-th adversarial token $x_{\text{adv}_{i}}$ can be denoted as:
$\begin{array}[]{c}x_{\text{adv}_{i}}=\mathop{\arg\min}\limits_{x_{\text{adv}_{i}}}\mathcal{L}.\end{array}$ (4)
To achieve this goal, TPIA propagates the loss value to the vocabulary embedding layer. Based on the gradient at token $x_{\text{adv}_{i}}$, it greedily searches for the suitable token $e^{\prime}_{i}$ in the vocabulary set of the Code LLM $V$ to replace $x_{\text{adv}_{i}}$, as shown in the following equation:
$e^{\prime}_{i}=\underset{e^{\prime}_{i}\in V}{\arg\min}\left(e^{\prime}_{i}-x_{\text{adv}_{i}}\right)^{T}\nabla_{x_{\text{adv}_{i}}}\mathcal{L},$ (5)
where $\nabla_{x_{\text{adv}_{i}}}Loss$ is the gradient of the task loss. TPIA uses beam search to enhance this token replacement strategy by considering the Top-$k$ token candidates (Line 6-7). This has been proven to be highly effective in existing NLP tasks [63, 50]. Afterwards, TPIA uses each token candidate to replace the original one to form a new $\bm{P^{\prime}}$ (Line 8-9). It then computes the new loss value using $\bm{P^{\prime}}$ and compares it with the original one: if the new value is better, $\bm{P}$ is replaced with $\bm{P^{\prime}}$ (Line 10-13). Finally, if $\bm{P}$ is not optimized within an entire cycle from 0 to $l$, TPIA considers it as optimal and terminates the optimization process (Line 16).
In particular, the design of our malicious perturbation generation is inspired by existing methods, such as universal trigger [50] and GCG [63]. We have already made some special designs, e.g., using Positon Code to locate the specific malicious target, based on characteristics of code completion and indirect prompt injection (instead of prompt injection). However, its effect is still very limited due to the difficulties of targeted malicious tasks. For example, as shown in Figure 9, its attack success rate (0-key, no FRE) is 0%. To address this, we propose Forward Reasoning Enhancement and Keyword-based Perturbation Design to further improve the perturbation generation process of our TPIA.
Figure 4: Loss function distortion during the forward reasoning process. When calculating the loss value using Equation 3, it optimizes the highest probability for generating the target token sequence. However, if the generation of the first target token fails, it can cause the loss function to become entirely inconsistent with the actual situation, making the perturbation ineffective in practical scenarios.
IV-D Forward Reasoning Enhancement
This optimization allows the optimization direction of adversarial computation to be closer to the actual reasoning process of Code LLMs (i.e., forward token sampling with randomness), thus improving the attack’s effectiveness and robustness. As shown in Figure 4, due to the operational mode of LLMs, where the final result is based on former sampling results, the original optimization function may deviate from actual conditions in certain situations. For example, the data deletion malicious instruction “os.system("rm -rf /")”, although brief, is difficult to integrate coherently with the surrounding context. This phenomenon is highly related to the optimization design in Equation (3), where TPIA calculates and sums the generation probabilities of individual tokens based on predetermined conditions. This indicates that a deviation occurring during the sampling of preceding tokens can distort the probability calculations for subsequent tokens.
To alleviate this problem, we propose a forward reasoning enhancement technique. Our intuition is that the former tokens of the target code should be assigned more weights to enhance the forward reasoning. Considering the computational overhead, we assign higher weights to the first $h$ tokens to mitigate the gradient calculation direction error. Its specific formulation is shown as follows:
$\mathcal{L}_{e}=\frac{1}{h}\sum_{i=1}^{h}\sum_{j=1}^{i}$ $\log(1-p(t_{i}|(\bm{T_{s}}\oplus\bm{P},t_{1},t_{2},\ldots,t_{j-1}))),$ (6)
$\mathcal{L}=(\mathcal{L}+r\mathcal{L}_{e})/(r+1),$
where $\mathcal{L}_{e}$ represents the additional reasoning-based enhancement applied to the original loss function, and $h$ is a hyperparameter set to either 1 or 2, which significantly improve the performance of attack in practice.
Figure 5: Examples of the keyword-based perturbation and other perturbations. The directly generated perturbation faces challenges in being effective while the manually-crafted perturbation lacks the stealthiness advantage. Our keyword-based perturbation can balance both effectiveness and stealthiness.
IV-E Keyword-based Perturbation Design
This optimization offloads insensitive parts of the malicious objective into static tokens, significantly reducing the total number of tokens required for a successful attack. From an empirical perspective, using fewer tokens and less obvious malicious characteristics will enhance the stealthiness of the attack. Compared to the manually-crafted perturbations, the illegibility of our malicious perturbations can prevent the victim from noticing their malicious intent. However, accurately conveying instructions using illegible tokens remains a significant challenge. A straightforward solution is to use a manually-designed perturbation as initialization, but this approach has two issues. First, the malicious perturbation cannot gain any advantage in terms of the token count. Second, this leads to ineffective optimization efforts. As shown in Figure 5, the manually-designed perturbation fails to not only reduce the token usage but also optimize out the core malicious features, leading to additional optimization rounds.
To alleviate these limitations, we propose another simple yet effective enhancement method, i.e., keyword-based perturbation design. Our idea is to transfer part of the information to static tokens, thereby reducing the information load on illegible tokens. Specifically, before each round of perturbation optimization in Algorithm 1, we use grid search to select the n tokens with the best performance as keywords. With the optimization step $s$, the equation can be formulated as follows:
$\bm{P}_{key}^{s+1}=\min_{\bm{P}_{key}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}\oplus\bm{P}_{adv}^{s},\bm{\theta}),$ (7)
$\bm{P}_{adv}^{s+1}=\min_{\bm{P}_{adv}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}^{s+1}\oplus\bm{P}_{adv},\bm{\theta}),$
where $\bm{P}_{key}$ and $\bm{P}_{adv}$ represents keywords part and adversarial part in perturbation $\bm{P}$ and the loss function is the same as Equation 6. To make the keywords relevant to the malicious objective, we select the tokens for the keywords from the target code and the position code, i.e., for token $p_{ki}$ in $\bm{P}_{key}$, there is $p_{ki}\in\{\text{\emph{target code}},\text{\emph{position code}}\}$. As shown in Figure 5, part of the target function name (TR) is the keywords, which is not directly related to the malicious objective (data deletion). By using a small number of statistical tokens, our method provides an effective design, maintaining the stealthiness and reducing the information load on ambiguous tokens.
V Experiments
In this section, we conduct an extensive evaluation of our TPIA, demonstrating its realistic threat to Code LLMs from the following three perspectives:
  1. 1)
    TPIA can successfully achieve representative malicious objectives on different open-source Code LLMs with high effectiveness and stealthiness. (Section V-B)
  2. 2)
    TPIA has a sufficient transferability to confirm its feasibility in real-world commercial applications under gray-box and black-box settings. (Section V-C)
  3. 3)
    TPIA’s modules are all effective in improving attack performance. (Section V-D)
V-A Main Settings

### v6

@@ 3.1 Formulation of External Prompt Injection Attacks
The Main Pipeline of Code Completion Task. Code completion is an essential task for Code LLMs, whose input may include code snippets, comments, pseudocode and user-specified natural language requests. Its output should consist of executable code blocks that meet user-specified requests. In this paper, we consider a general task format. Specifically, the input code block is defined as $\bm{X}=\{x_{1},x_{2},\ldots,x_{m}\}$, where $x_{i}\in\bm{X}$ can be any token, and $\bm{X}$ as a whole adheres to the code format requirements. For instance, natural language requests can only be included in the code comments. The output is the completed code block $\bm{Y}=\{y_{1},y_{2},\ldots y_{n}\}$, which should be logically executable and functionally compliant with the requirements. We denote a Code LLM with weights $\bm{\theta}$ as $\Theta$, and the code completion task is simplified as $\bm{Y}=\Theta(\bm{X})$.
The Formulation of EPI. In general, there are three parties, including a Code LLM-integrated application, a user, and an attacker, involved in our EPI. The user employs the Code LLM-integrated application to complete the code project. The Code LLM-integrated application constructs the input $\bm{X}$ for the integrated Code LLM based on the code project and get the output. The attacker generates a concise induced perturbation $\bm{P}$ and introduces it into the input $\bm{X}$. The Code LLM, affected by $P$, will generate code that is functionally close to the attacker’s malicious objective $O$, i.e.,
$\mathcal{D}(O,\mathcal{F}(\Theta(\bm{X}\leftarrow\bm{P})))<\epsilon,$ (1)
where $\mathcal{F}(\cdot)$ denotes the functionality of the code, $\bm{X}\leftarrow\bm{P}$ means injecting $\bm{P}$ into $\bm{X}$. $\mathcal{D}(\cdot,\cdot)$ is a function that measures functional differences. When its value is less than $\epsilon$, it indicates functional equivalence between inputs.
Attacker’s Capacities and Attack Scenarios. We assume that the attacker can obtain the parameters of an surrogate Code LLM to conduct attacks. This assumption is consistent with many adversarial methods [26, 51]. Specifically, we consider three main attack settings: white-box, gray-box, and black-box settings. The former assumes that the attacker knows the specific parameters of Code LLM, which generally corresponds to scenarios where the model is an open-sourced model. The latter two assume that the attacker only knows the model series (rather than the specific version) and knows nothing about it, respectively, which generally corresponds to scenarios of using commercial Code LLM-intergrated applications. In our attack scenarios, the attackers introduce external code to the victim user’s code project. We argue that it can be easily and stealthily achieved in the real world. For example, introducing external code by copying or installing packages is commonly done during the programming process [23, 18]. Meanwhile, external code from open-source platforms (e.g., GitHub) is usually downloaded as a knowledge base via retrieval-augmented generation [22], which can also be used as input to improve the performance of Code LLMs [21]. In particular, to avoid detection by existing methods (e.g., static analysis [24] and sandboxing [25]), external code should not immediately contain malicious functionalities, i.e., attackers should use non-functional code as induced perturbation.
3.2 Overview
In this section we briefly introduce our greedy perturbation optimization method (dubbed ‘ShadowCode’) for automatic and effective external prompt injection. As illustrated in Figure 2, ShadowCode operates in two sequential stages: (1) design optimization objectives and (2) perturbation optimization. In the first stage, we design the perturbation optimization objectives for induced perturbation based on the attacker’s malicious objective. In the second stage, we employ the greedy search to optimize the induced perturbation.
3.3 Design Optimization Objeciteves
In this stage, we first describe how to convert diverse malicious objectives to structured output tuples. Then we expand the output tuple to a full code contextual simulation. Finally, we design a loss function that leverages the simulation to guide the search for the optimal perturbation.
Output Tuple. We hereby describe how to convert the malicious objective to an output tuple. Let the normal code completed by a Code LLM is denoted by $\bm{Y}$, the implementation of malicious purposes is maliciously modifying $\bm{Y}$. Such modifications can be divided into three types: deletion, addition, and alteration, all of which can be represented by an output tuple (position code $\bm{Y}_{P}$, target code $\bm{Y}_{T}$). Specifically, for deletion, the attacker wants to delete the code segment $[y_{i},y_{i+1},\ldots,y_{i+k}]$ in $\bm{Y}$, and we have $\bm{Y}_{P}=[y_{p},y_{p+1},\ldots,y_{i-1}]$ and $\bm{Y}_{T}=[y_{i+k+1},y_{i+k+2},\ldots,y_{t}]$; For addition, the attacker wants to add the code segment $\bm{T}$ into the position between $y_{i}$ and $y_{i+1}$, and we have $\bm{Y}_{P}=[y_{p},y_{p+1},\ldots,y_{i-1}]$ and $\bm{Y}_{T}=[\bm{T},y_{i+1},y_{i+2},\ldots,y_{t}]$; For alteration, the attacker wants to change the code segment $[y_{i},y_{i+1},\ldots,y_{i+k}]$ to $\bm{T}$, and we have $\bm{Y}_{P}=[y_{p},y_{p+1},\ldots,y_{i-1}]$, $\bm{Y}_{T}=[\bm{T},y_{i+k+1},y_{i+k+2},\ldots,y_{t}]$.
Code Contextual Simulation. Optimizing the induced perturbation based solely on the output tuple ignores the surrounding code context. To ensure the effectiveness of ShadowCode in uncontrolled context environment, we expand the output tuple to simulate a code context of the victim user’s porject. In general, although the contextual environment of the victim user’s project is usually uncontrollable for the attacker, there are still some fundamental logics to follow for code projects and code generation processes, which can be exploited by ShadowCode to simulate a realistic environment. Accordingly, our code context simulation is constructed based on the following three logics:
@@ 3.3 Design Optimization Objeciteves
Logic 1: Necessary Condition Argument. The code environment configured by the user must meet the basic requirements for the malicious objective, such as the necessary packages and method declarations. Therefore, we introduce this kind of code as conditional code at the beginning of the whole simulated code.
Logic 2: Fixed Generation Order. Code generation usually proceeds in a forward textual order. Accordingly, the target code should be inserted immediately after the position code to achieve the malicious objective and the subsequent content of the target code is not strongly relevant to the objective. Thus, we connect the position code and target code and place them at the end of the simulated code.
Logic 3: Noise in Context. Since realistic projects usually consist of code with different functions, the code context is likely to contain noise code not entirely related to malicious purposes, which may have a potential impact on the effectiveness of the final induced perturbation. As such, we design the noise code, adding irrelevant noise to enhance the robustness of the induced perturbation.
@@
Based on three logics above, we can construct the code contextual simulation for different malicious objectives. To better illustrate our strategy, we provide a straightforward example in Figure 3, where the malicious objective is to make the victim use a weak encryption mode. The task components are automatically generated with the following stages. First, the position code should contain a fundamental encryption situation: the encryption key “key”, the text to be encrypted “iv” and the first half of the encryption instruction “cipher = DES.new(key,”. These variables and functions are necessary for the encryption methods, e.g., “CBC”. To achieve the malicious objective, the complete code instruction should call the weaker encryption method, “ECB”, thus the target code is “DES.MODE_ECB)”. After obtaining the output tuple, ShadowCode then declares the Crypto series methods that enable the encryption code to run normally in the conditional code. Finally, ShadowCode constructs the code that is not related to the task and uses it as the noise code. The final code contextual simulation is constructed from multiple aspects for the same task to collaboratively ensure the effectiveness.
Figure 3: The example of code contextual simulation, where the attacker aims to induce the Code LLM to use encryption mode with low security (i.e., ECB encryption). The simulation consists of four code segments: conditional code, noise code, position code, and target code, to simulate the context environment most relevant to the malicious objective.
Loss Function of Perturbation Optimization. After obtaining the code contextual simulation, a loss function can be formulated for optimizing induced perturbations. To achieve the malicious objective, the optimization aims to identify the perturbation that maximizes the probability of realizing the simulated successful attack. Specifically, the simulated contextual code is first divided into input and output components. The input to the Code LLM comprises conditional code, noise code, and position code, simulating a code completion task that can be exploited to achieve the malicious objective. This entire input is referred to as task code. The output, termed target code, serves as the label for the perturbation optimization task. Subsequently, with $\bm{Y}_{T}=[t_{1},t_{2},\ldots t_{k}]$, the loss function is defined as:
$\mathcal{L}=$ $\sum_{i=1}^{|\bm{Y}_{T}|}\log(1-\mathbb{P}(t_{i}|\bm{T}_{s}\leftarrow\bm{P},t_{1},t_{2},\ldots,t_{i-1})),$ (2)
where $\bm{T}_{s}$ represents the task code, the probability $\mathbb{P}$ is inferred and calculated using the surrogate model weights $\bm{\theta}$. This function aims to maximize the likelihood of generating the target code given the task code as input. When the Code LLM receives the perturbation, it is induced to generate the target code if the context is appropriate (i.e., similar to the simulation in code functionality), thereby achieving the attacker’s malicious objectives.
3.4 Perturbation Optimization
In this stage, we first introduce two enhancement modules for the optimization process: forward reasoning enhancement and keyword-based perturbation design. Then, we will describe the whole algorithm for induced perturbation optimization.
Forward Reasoning Enhancement. Directly employing the loss function defined in Equation (2) may lead to distortions during the autoregressive generation process. As shown in Figure 4, since token prediction results are based on previous sampling outcomes, the original loss function may deviate from actual conditions due to unexpected predictions of the first few tokens. For example, the data deletion malicious instruction “os.system("rm -rf /")”, although brief, is difficult to integrate coherently with the surrounding context (i.e., immediately after calling a demo function “TRG”). If the induced Code LLM fails to generate the first token “os”, the probabilities of generating subsequent tokens will diverge significantly.
Figure 4: Distortion during the autoregressive generation process. The attacker aims to insert a malicious deletion command after the position code, whereas the Code LLMs would typically generate the result printing command. During the perturbation optimization process, the generation probability of the first token in the target code, i.e., “os”, is not adequately optimized. Consequently, the Code LLM may not be successfully induced upon encountering the position code and generate an undesired token “print”. Subsequently, the failure to generate the first token compromises the successful generation of subsequent tokens.
To alleviate this problem, a forward reasoning enhancement module is introduced. Motivated by the aforementioned discussion that subsequent tokens are influenced by preceding tokens in an autoregressive generation manner, earlier tokens hold greater importance for the overall generation process of the target code. Therefore, the importance of the initial tokens for the loss value is increased by assigning them higher weights, thereby enabling the optimization objective to adapt to the autoregressive mechanism. This increase in weight can be achieved through adding an additional loss. Its specific formulation is shown as follows:
$\mathcal{L}_{e}=\frac{1}{h}\sum_{i=1}^{h}\sum_{j=1}^{i}$ $\log(1-\mathbb{P}(t_{i}|(\bm{T_{s}}\leftarrow\bm{P},t_{1},t_{2},\ldots,t_{j-1}))),$ (3)
$\mathcal{L}=(\mathcal{L}+r\mathcal{L}_{e})/(r+1),$
where $\mathcal{L}_{e}$ is the additional loss, $r$ is its coefficient, and $h$ is the number of enhanced tokens.
Figure 5: Examples of different induced perturbations. The attacker aims to induce the Code LLM to generate a data deletion command (“os.system("rm -rf /")”) immediately after a specified function call (“TRG”) via induced perturbation. Manually designed induced perturbations lack stealthiness due to their obvious malicious intent, a flaw that a simple greedy search-based replacement alone cannot rectify. Induced perturbations optimized with random initialization necessitate extensive token usage to ensure effectiveness. Our method combines optimized induced perturbations with a prefix (keyword), thereby achieving both effectiveness and stealthiness with limited tokens.
Keyword-based Perturbation Design. This module aims to significantly reduce the token count required for induced perturbations by offloading information into prefix tokens, which enhances attack stealth. Arguably, minimizing token usage and obscuring malicious characteristics are critical to maintaining attack stealth, thereby avoiding detection by victim users or RAG’s developers. Generally, perturbations optimized via greedy search typically exhibit less obvious malicious characteristics compared to manually designed ones. However, these perturbations struggle to accurately convey malicious instructions as effectively as manually designed perturbations, resulting in higher token usage. A seemingly straightforward solution, which involves optimizing based on manually designed perturbations by replacing obvious malicious characteristics with equivalent tokens, proves difficult in practice. As shown in Figure 5, the tokens constituting these obvious malicious characteristics (i.e., convey dangerous data deletion instructions) are the most critical and challenging to replace via greedy search.
To alleviate these limitations, we propose a simple yet effective enhancement method, i.e., keyword-based perturbation design. Intuitively, the purpose of the induced perturbation is to fulfill the attacker’s malicious objective. Therefore, the output tuple, which expresses malicious objective, will contain the necessary information for the induced perturbation. Motivated by this understanding, we leverage tokens from the output tuple to construct a prefix for the perturbation, termed a keyword, which is then injected into the victim’s code context alongside the perturbation. During keyword construction, we use the loss function in Equation (3) to greedily select the optimal token combination. In particular, we utilize only a very small number of tokens (typically one or two) for keyword construction to ensure no overt malicious intent is present in the injected content. As shown in Figure 5, the keyword “TR” exhibits no overt malicious intent but can enhance the perturbation’s effectiveness compared to directly optimized counterparts.
Optimization Algorithm. We hereby introduce the specific optimization process of ShadowCode for induced perturbations and their corresponding keywords. Specifically, we employ greedy gradient search to optimize an initialized perturbation $\bm{P}=[p_{1},p_{2},\ldots p_{l}]$, where $l$ denotes the token count. As shown in Algorithm 1, our ShadowCode method optimizes the perturbation token by token in each iteration (line 4). We utilize the loss function in Equation (3) for greedy gradient search (line 5), i.e., the optimization objective for the $i$-th token $p_{i}$ is:
$\begin{array}[]{c}p_{i}=\mathop{\arg\min}\limits_{p_{i}}\mathcal{L}.\end{array}$ (4)
To achieve this, ShadowCode propagates the loss value to the vocabulary embedding layer. Based on the gradient at token $p_{i}$, it greedily searches for the suitable token $e^{\prime}_{i}$ within the Code LLM’s vocabulary set $V$ to replace $p_{i}$, as shown in the following equation:
$e^{\prime}_{i}=\underset{e^{\prime}_{i}\in V}{\arg\min}\left(e^{\prime}_{i}-p_{i}\right)^{T}\nabla_{p_{i}}\mathcal{L},$ (5)
where $\nabla_{p_{i}}\mathcal{L}$ represents the gradient of the task loss. ShadowCode enhances this token replacement strategy by employing beam search, which considers the Top-$k$ token candidates (Line 6-7). ShadowCode then replaces the original token with each candidate to form a new $\bm{P^{\prime}}$ (Line 8-9), and computes a new loss value $\mathcal{L}^{\prime}$: if $\mathcal{L}^{\prime}$ is better than $\mathcal{L}$, $\bm{P}$ is replaced with $\bm{P^{\prime}}$ (Line 10-13). Finally, if $\bm{P}$ is not optimized within an entire iteration, ShadowCode considers it optimal and terminates the optimization process (Line 16).
Algorithm 1 Induced Perturbation Optimization
@@ 3.4 Perturbation Optimization
1:  Input: Task Code $\bm{T_{s}}$, target code $\bm{Y}_{T}$, initial perturbation $\bm{P}$ with tokens $p_{1:l}$, beam search size $k$
2:  Vocab set $V$, weights $\bm{\theta}$ of the surrogate Code LLM
@@
5:        $\mathcal{L}=\text{get\_loss}(\bm{T_{s}},\bm{Y}_{T},\bm{P},\bm{\theta})$
6:        $\mathcal{G}=\nabla_{p_{i}}\mathcal{L}$
7:        $\mathcal{P}_{\text{candidates}}=\text{Top-}k_{e^{\prime}_{i}}\text{min}({e^{\prime}_{i}}^{T}\mathcal{G}),e^{\prime}_{i}\in V$
@@
9:           $\bm{P}^{\prime}:\text{replace }p_{i}\text{ to }\mathcal{P}_{\text{candidates}}^{j}$
10:           $\mathcal{L}^{\prime}=\text{get\_loss}(\bm{T_{s}},\bm{Y}_{T},\bm{P^{\prime}},\bm{\theta})$
@@
16:  until $\bm{P}$ does not change in the whole iteration
@@
Additionally, we use grid search to optimize the keyword. The perturbation and the keyword are optimized alternately, which can be expressed as follows:
$\bm{P}_{key}^{s+1}=\arg\min_{\bm{P}_{key}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}\oplus\bm{P}^{s},\bm{\theta}),$ (6)
$\bm{P}^{s+1}=\arg\min_{\bm{P}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}^{s+1}\oplus\bm{P},\bm{\theta}),$
where $s$ is the number of optimization steps and $\bm{P}_{key}$ represents the keyword. $\oplus$ means concatenating two token sequences. We select keyword tokens from the output tuple, i.e., $\forall p_{ki}$ in $\bm{P}_{key},p_{ki}\in\{\text{\emph{target code},},\text{\emph{position code}}\}$. For example, as shown in Figure 5, the keyword “TR” are tokens from the position code of the output tuple.
4 Experiments
4.1 Main Settings